Compare 3 groups with 21 observations per group, MSE 50, and MSG 340. Fill in the blanks.

DFG =

DFE=

F=
Answer #1

number of groups k = 3

Number of observation n = 3 * 21 = 63

DFG = k - 1 = 3 - 1 = 2

DFE = n - k = 63 - 3 = 60

F = MSG / MSE = 340 / 50 = 6.8